Calculate the density of nitrogen gas at 120 degrees Celsius and a pressure of 700mmHg
1
Expert's answer
2020-04-24T13:34:48-0400

T = 120 0C = 273+120 K = 393 K

P = 700 mmHg = 101325/760*700 Pa = 93326 Pa

M(N2) = 28 g/mol

According to Ideal gas law,

"PV = nRT\\
\rho = {\frac {m}{V}}={\frac {nM}{V}}\\
P={\frac {n}{V}}RT\\
{\frac {n}{V}}={\frac {P}{RT}}\\
\rho = {\frac {P}{RT}}M={\frac {93326}{8.314*393}}*28 =800" g/m3 = 0.8 g/L